System Status Archive

This page provides current and archived status updates for Other


View Systems Overview
Current Status
OtherPower Outage in Olin, Chemistry, BiologyThere is no electricity to Olin Science, Rooke Chemistry and Biology buildings. As a result, technology in those areas is also unavailable.Posted: Wednesday, October 1, 2014 at 8:20 AM


Archived Status Alerts for Other


 TitleDescriptionSubmittedResolved
Voicemail UnavailableThe voicemail system is currently unavailable. We are working to resolve the problem.10/1/14 08:48 10/1/14 12:07
Voicemail UnavailableThe voicemail system is currently unavailable. We are working to resolve the problem.10/1/14 09:51 10/1/14 09:52
ARTSTOR and Shared Shelf OutageThe ARTSTOR database and Shared Shelf service will be down for maintenance from 5-6pm on Wednesday 9/24. If you have questions, please contact library@bucknell.edu9/24/14 16:14 9/24/14 18:15
Bmail/Google DriveGoogle Drive appears to be intermittently unavailable at this time. 9/16/14 09:28 9/16/14 15:49
Tedy outageWe are currently experiencing performance issues with Tedy (Team Dynamix) and are working to resolve the issue. There will be a brief outage at 12:15. The services will be unavailable for 5 minutes. f you have any questions or concerns please feel free to contact Tech Support at techdesk@bucknell.edu or by calling 570-577-7777.9/15/14 11:14 9/15/14 12:20
EDGE Scheduled MaintenanceEDGE will be unavailable due to scheduled maintenance from 8:30 pm on Friday, September 12 until 1:30 am on Saturday, September 13. During this outage performance enhancements and new features will be enabled. EDGE is the faculty and staff learning, development, and performance system.9/9/14 08:49 9/13/14 10:00
B-bill is currently unavailableB-bill is currently unavailable. The University is working with the vendor to solve the issue and hope to have it resolved quickly.9/11/14 08:37 9/11/14 09:00
Mozy Pro Intermittent AccessMozy Pro is currently experiencing intermittent issues on their servers that may affect your ability to backup. This does not affect any data currently backed up to the Mozy Pro servers.9/4/14 13:54 9/4/14 15:31
EPM 11 - reports not completingSome reports are not running to completion. May get a Server [2009]; Can't allocate resources or reports hanging and not completing. We have a work-around pending testing.8/28/14 10:47 9/2/14 13:00
EPM 11 not availableEPM 11 is currently unavailable. We are working to bring it back as soon as possible.8/30/14 10:31 8/30/14 10:47
Concur Intermittent Login ProblemsConcur users may experience intermittent problems logging in. If you are able to login, please do not logout. We are working with the vendor to resolve the problem.8/28/14 09:29 8/29/14 14:53
Concur UnavailableConcur Travel and Expense (B-Travel) is currently unavailable due to vendor issues. We will update this message when the problem has been resolved.8/27/14 14:08 8/27/14 16:27
Campus Network Service OutageDear Campus Community: There will be an on-line service outage affecting campus and downtown offices beginning at 8:00 P.M. on Friday, August 8 and continuing until Noon on Sunday, August 10. During this time the only services that will be available will be Bmail (and related applications such as Calendar, Drive, etc.) and the outward-facing Bucknell website. All other network services (Banner, Netspace, myBucknell and associated services like jobs@bucknell, B-bill, etc.) will be inaccessible. We expect campus telephone service will also be unavailable with the exception of emergency phones and cellular service, which will continue to operate. During the outage you may still reach Public Safety at 570-577-3333 and the Bucknell Switchboard at 570-577-2000. The campus network, both wired and wireless, will be available during the maintenance period allowing access to the Internet. More details about this project may be found by visiting: http://lit.blogs.bucknell.edu/?p=6864 We greatly appreciate your understanding. If you have any questions or concerns please feel free to contact Tech Support at techdesk@bucknell.edu or by calling 570-577-7777. 8/7/14 10:25 8/10/14 13:30
Research by Subject Outage, Tuesday August 5Due to a systems upgrade, the Research by Subject pages will potentially be unavailable from 6:00am until noon on Tuesday August 5. If you need to access research databases during the outage, you can do so directly from the A-Z Database list at: http://my.bucknell.edu/script/LibraryIT/Databases/ Questions or concerns may be directed to the Library Services Desk at 570.577.14628/4/14 11:12 8/5/14 12:00
Telephone Outage Early Saturday Dear Campus Community: There is a brief campus telephone outage scheduled from 6:00 A.M. to 8:00 A.M. on Saturday, August 2. Network upgrades related to the telephone system will be performed during this time. Cellular service is not affected by this outage. Thank you for your understanding during this scheduled maintenance event. 8/1/14 10:02 8/2/14 08:30
Online Learning Center Scheduled MaintenanceThe Online Learning Center (training.bucknell.edu) will be unavailable due to scheduled maintenance for up to 20 minutes on Saturday, June 28 between 1:00 am and 3:00 am. Users who attempt to login may receive a message indicating that the system is undergoing system maintenance.6/25/14 13:15 6/30/14 08:00
University Switchboard UnavailableThe Switchboard is experiencing technical difficulties and is currently unable to receive calls. If needed, please call Public Safety at 7-3333.6/24/14 09:04 6/24/14 09:35
On-Line Learning Center Currently UnavailableDear Faculty and Staff: The Bucknell On-Line Learning Center is currently unavailable. Our staff is looking into the cause of the problem. Thank you for your patience.6/17/14 12:49 6/18/14 08:23
Bmail DriveDrive is not loading - there appears to be issues on Google's end. 6/4/14 10:37 6/9/14 13:22
Self Service for Macs Briefly Unavailable Tomorrow MorningDear Faculty and Staff: For those of you with Macintosh computers, Self Service will be unavailable from 8:15 A.M. to 8:30 A.M. on Thursday, May 1 while maintenance is performed. Thank you for your patience during this scheduled service activity.4/30/14 13:51 5/1/14 09:00
TeamDynamix (Tedy) inaccessible 4/16 from 12:15-12:30In order to complete maintenance on the Team Dynamix server the system will be inaccessible from 12:15 to 12:30 on Wednesday April 16th.4/16/14 11:27 4/16/14 12:22
EPM 11/Hyperion currently not available through ASAPThere has been an issues with accessing EPM11/Hyperion through the ASAP launch page after the access to it was removed from other pages. 4/14/14 09:21 4/14/14 09:48
BUTV missing some channelsSome channels are showing only a green screen this morning. We're working with Windstream, our provider, to determine the source of the problem. Update: 10:30am (from Windstream) WBRE, WYOU and WOLF are off air due to loss of power on the mountain. PP&L Crews are on Penobscot to replace a span of power line that went down due to EXTREMELY SEVERE ICING. Current estimate for restoration is 11:30 EDT. This is an estimate.4/1/14 10:24 4/1/14 12:03
Notes...
Close

Places I've Been

The following links are virtual breadcrumbs marking the 27 most recent pages you have visited in Bucknell.edu. If you want to remember a specific page forever click the pin in the top right corner and we will be sure not to replace it. Close this message.